whats the probability of getting a 6 or 8 on two dice? whats the probability of Not getting a 6 or 8?
Found 2 solutions by Nate, sowjanya:
Answer by Nate(3500)

Total Chances: 36
Prob of Six:
Die: 1,5
Die: 5,1
Die: 2,3
Die: 3,2
Chances: 4
4 / 36 = 1 / 9 or about 11.1%
Prob of Eight:
Die: 2,6
Die: 6,2
Die: 3,5
Die: 5,3
Die: 4,4
Chances: 5
5 / 36 or about 13.9%
Not getting either:
100% - 11.1% - 13.9% = about 75%

Answer by sowjanya(1)

Total Chances: 36
Prob of getting Six:
Die: 1,5
Die: 5,1
Die: 3,3
Die: 2,4
Die:4,2
Chances: 5
5/ 36 Prob of Eight:
Die: 2,6
Die: 6,2
Die: 3,5
Die: 5,3
Die: 4,4
Chances: 5
5 / 36
Probality of not getting either:
1-(5/36)-(5/36)
=26/36
=13/18
=72%